Please Help!!

Your firm just finished the year, in which it had cash earnings of $400. You forecast your firm to have a quick growth phase from year 0 to year 5, in which it grows at a rate of 40% per annum. Your firm’s growth then slows down to 20% per annum between year 5 to year 10. Finally, beginning in year 11, you expect the firm to settle into its long-term annual growth rate of 2% . You also expect your cost of capital to be 15% over the first 5 years, then 10% over the next 5 years, and 8% thereafter. What do you think your firm is worth today? (Advice: Use a computer spreadsheet program.)

My answer was $41,387 just making sure this is right!
Answer #1
Year Earning Cost of capital Present Value (PV)
A 0 $400 15% $400.00 PV at year 0
B=A*1.4 1 $560 15% $486.96 PV at year 0
C=B*1.4 2 $784 15% $592.82 PV at year 0
D=C*1.4 3 $1,098 15% $721.69 PV at year 0
E=D*1.4 4 $1,537 15% $878.58 PV at year 0
F=E*1.4 5 $2,151 15% $1,069.57 PV at year 0
G=F*1.2 6 $2,581.56 10% $2,346.87 PV at Year 5
H=G*1.2 7 $3,097.87 10% $2,560.22 PV at Year 5
I=H*1.2 8 $3,717.44 10% $2,792.97 PV at Year 5
J=I*1.2 9 $4,460.93 10% $3,046.87 PV at Year 5
K=J*1.2 10 $5,353.11 10% $3,323.86 PV at Year 5
L=K*1.02 11 $5,460.18 8%
PV at Year 10 for Earning from Year 11 to perpetuity $91003 (5460.18/(0.08-0.02))
PV at Year 5 for Earning from Year 11 to perpetuity $56505.70 (91003/(1.1^5)
A PV at Year 0 for Earning from Year 11 to perpetuity $28093.32 (56505.70/(1.15^5)
PV at year 5 for earning of year6 to10 $14,070.79 ((2346.87+2560.22+2792.97+3046.87+3323.86)
B PV at year 0 for earning of year6 to10 $6,995.67 (14070.79/(1.15^5)
C PV at year 0 for earning of year0 to 5 $4,149.62 (400+486.96+592.82+721.69+878.58+1069.57)
W=A+B+C Firm's Worth Today $39238.61
Answer should be $39238.61